OpenStudy (anonymous):

Please help! Will give Medal!! A cone has a volume of 5 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? 15 in3 20 in3 25 in3 30 in3

OpenStudy (anonymous):

@sammixboo @Zarkon @StudyGurl14 @genny7

OpenStudy (campbell_st):

let the cone and cylinder both have radius r and height h the volume of the cone \[V = \frac{1}{3} \pi r^2h\] the cylinder \[V = \pi r^2 h\] so what do you think..?

OpenStudy (campbell_st):

how many cones fit in the cylinder...? use the same rule for your answer

OpenStudy (anonymous):

i think the volume of the cylinder is a third of 5.... right?